one number is 6 more than a smaller number. if the difference between 2 times the large number and 5 times the smaller number is 3. find both numbers.
Answer by Cromlix(4381)

Hi there,
Make the smaller number = 'x'
Other number = x + 6
The difference between 2 times
the larger number and 5 times
the smaller is 3
2(x + 6) - 5(x) = 3
2x + 12 - 5x = 3
Collect like terms:
2x - 5x = 3 - 12
- 3x = -9
x = 3
Smaller number = 3
Other number = 9
Hope this helps :-)